The number of errors in Abel’s presentations has a mean of 2 per page. The moderator randomly selects one page, what is the probability that a) there is no error and b) there are 5 errors
Answer #1

X ~ Poisson ( )

Where mean = 2

P(X) = e-X / X!

a)

P (X = 0) = e-2 20 / 0!

= 0.1353

b)

P (X = 5) = e-2 25 / 5!

= 0.0361